A radioactive sample of thallium-207 has a half-life of 4.77 minutes has an activity of 1500 Ci.
What is the mass of the sample of thallium? (to 2 s.f and in µg)
1
Expert's answer
2020-05-29T10:03:35-0400

From the formula for the activity of a substance

"C=\frac{0.693 \cdot N_A \cdot m}{M \cdot T_{1/2}}"

Where

"N_A=6.022 \cdot 10^{26}" - Avogadro number

"M" - atomic mass of the radioactive element

"T_{1/2}=4.77 \cdot 60=286.2 s" - half-life of a radioactive element

"m" -mass of a radioactive element

Note

"m=\frac{C \cdot M \cdot T_{1/2}}{0.693 \cdot N_A}"

Given that

"1Ci=3.7 \cdot 10^{10}{decays}/s"

"m=\frac{1500 \cdot 206.977 \cdot( 4.77\cdot 60 )\cdot 3.7 \cdot 10^{10} \cdot 10^{6}}{0.693 \cdot 6.022 \cdot 10^{26}}=7.878 \cdot 10^{-3}\mu g"